OpenStudy (anonymous):

Solve this equation for A: A ÷ 2 = 4

OpenStudy (lgbasallote):

A/2 = 4 cross myltiply A = 8

Directrix (directrix):

A divided by 2 = 4 A/2 = 4 Multiply both sides by the reciprocal of (1/2) 2(A/2) = 2(4) A = 8